An op-amp with slew rate 6 V/μs and gain of 20 dB. If this amplifier must faithfully amplify sinusoidal signal from dc to 120 KHz without introducing any slew rate induced distortion, then the input signal level must not exceed


1
Expert's answer
2021-09-03T01:59:58-0400

If an op amp is operated above its slew rate limit, signals will become distorted.

Given that:

Slew Rate: 6 V/μs

Gain of 20 dB

From dc to 120 KHz

The decibel (dB) value is a ratio used for comparing and calculating levels of change in power

and is not the power itself.

Then input level is given by:


dB = 10log 10 [P 2 /P 1 ]



If P 2 /P 1 is equal to 1, that is P 1 = P 2 then:

Let x be input signal level:


dB = 10log 10 [1] = log 10 [x/10] = log 10 [0.1] = antilog[0.1]

20= 10log 10 [1] = log 10 [x/10] = log 10 [0.1] = antilog[0.1]


We get the value of: 10 0.1 = 1.259